OpenStudy (prettygirl_shynice):

wil medel + fan which proportion can be used to determine what number is 15% of 30? a.15/30=x/100 b.15/100=x/30 c.15/100=30/x d.30/100=15/x help and explain how i do it plz help

OpenStudy (mathmale):

You might actually do the work to find 15% of 30, and then look backwards to determine which of the four possible answers is closest to what you've actually done.

OpenStudy (prettygirl_shynice):

i dont know how to do it

OpenStudy (mathmale):

Find 15% of 30. 1. Convert 15% to a decimal fraction: 15/100 = 0.15 2. Multiply 30 by 0.15 (could you do this, please?)

OpenStudy (prettygirl_shynice):

yes

OpenStudy (prettygirl_shynice):

okay its 4.5

OpenStudy (mathmale):

that's right! that 4.5 is 15% of 30. Now, go back and review the four answer choices. Which one shows an approach closest to the one we've used here?

OpenStudy (prettygirl_shynice):

ok

OpenStudy (prettygirl_shynice):

omm b

OpenStudy (mathmale):

Let's check out (b). b.15/100=x/30 Substitute x=4.5. (We saw earlier that 15% of 30 is 4.5.)

OpenStudy (prettygirl_shynice):

yep

OpenStudy (mathmale):

\[\frac{ 15 }{ 100 }=?=\frac{ 4.5 }{ 30 }\]

OpenStudy (mathmale):

I'm asking you to determine whether or not this equation is true.

OpenStudy (prettygirl_shynice):

omm true?

OpenStudy (usukidoll):

try using cross multiplication (15)(30) = (4.5)(100)

OpenStudy (mathmale):

Yes, you'll need to prove mathematically whether or not the equation is true. Please cross-multiply, as U. D. has suggested.

OpenStudy (prettygirl_shynice):

ok iso multiply 15 time 100 and 30 time 4.5 ?

OpenStudy (prettygirl_shynice):

ment 15 time 30 and 100 time 4.5

OpenStudy (mathmale):

Actually, cross-mult in the case of \[\frac{ 15 }{ 100 }=\frac{ 4.5 }{ 30 }\]... would be appropriate. 15*30 is one such mult.; the other is 4.5*100. Determine wehther is this is true or false, please.

OpenStudy (prettygirl_shynice):

true

OpenStudy (mathmale):

Way cool. That means that your choice of (b) was appropriate!

OpenStudy (mathmale):

Always check your results, whenever that's possible.

OpenStudy (prettygirl_shynice):

so im wrong

OpenStudy (mathmale):

Why would you think that? You found that (15)(30)=450 and that 4.5(100)=450. Is 450=450?

OpenStudy (prettygirl_shynice):

yea

OpenStudy (mathmale):

that means your answer is correct!

OpenStudy (prettygirl_shynice):

oh im tag u in my next question and thx